OpenStudy (anonymous):

Can someone please help solve this differential equation: dy/dx=[x(y^2-1)]/[2(x-2)(x-1)]

OpenStudy (anonymous):

Looks separable.

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

\[\implies \frac{1}{y^2 - 1}\, dy = \frac{x}{2(x-2)(x-1)}\,dx\]

OpenStudy (anonymous):

I've gotten it that far, it's the integrals at that point that I'm having trouble with :/

OpenStudy (anonymous):

Yeah, they don't look particularly fun, but doable. The first one looks like a trig or hyperbolic trig function.

OpenStudy (anonymous):

For the second, I'd look for a partial fraction decomposition.

OpenStudy (anonymous):

thanks so much! i'll try that!

OpenStudy (anonymous):

lhs = -tanh x

OpenStudy (anonymous):

is the right hand side ln(x-2)-(ln(x-1)/2)?